OpenStudy (anonymous):

Solve the DE (x^2-2x+3) y''+18xy'-9xy=0 using power series

OpenStudy (perl):

what class is this?

OpenStudy (anonymous):

differential equations

OpenStudy (anonymous):

Let \[y=\sum_{n=0}^{\infty} a_n x^n\] Sub into DE: \[=\sum_{n=2}^{\infty}n(n-1)a_n x^n -2\sum_{n=1}^{\infty}n(n+1)a_{n+1}x^n+3\sum_{n=0}^{\infty}(n+2)(n+1)a_{n+2}x^n\] \[+18\sum_{n=1}^{\infty}na_nx^n-9\sum_{n=1}^{\infty}a_{n-1}x^n=0\] This is where I am stuck! Any help?

OpenStudy (anonymous):

About \(x=0\), I presume?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

any help?